OpenStudy (anonymous):

A sample of gas occupying 350. ml at 25°C is cooled to -25°C. What volume will it occupy if the pressure remains constant?

OpenStudy (zale101):

PV=nRT Based on your question what's changing? I only see temperature and volume, the rest is constant So, eliminate moles, pressure, and since R is always constant we'll just remove it anyways. V=T <--only left with this \(\Large \frac{V_1}{T_1}=\frac{V_2}{T_2}\) v1=\(350. ml*\frac{1L}{1000ml}=.35~L\) t1=\( 25°C +273K=298~K\) v2= Unknown t2=\(-25°C +273K=248~K\) \(\Large \frac{.35~L}{298~K}=\frac{V_2}{248~K}\) Solve for v2